OpenStudy (anonymous):

What is the equation of a line that is parallel to y = 3x – 6 and passes through the point (-8,2) Write it in slope-intercept form.

OpenStudy (littlebird):

Example of slope intercept formula in the image.

OpenStudy (anonymous):

y= -4 + 6 ?

OpenStudy (anonymous):

@saify_12

OpenStudy (littlebird):

I made a typo in my calculations. Im fixing it now.

OpenStudy (littlebird):

y=3x+20

OpenStudy (littlebird):

Parallel lines have the same slope, but a different y- intercept. Plug in the values from the given point and leave the y-intercept as an unknown variable(b). y=3x-6 (2)=3(-8)+b 2=-18+b b=20 Now write your final equation.